Which of the following investments will produce the largest amount of interest?

a. $1,000 at 6% interest compounded annually for 5 years
b. $1,000 at 7% interest compounded annually for 4 years
c. $1,000 at 7% interest compounded quarterly for 4 years
d. $1,100 at 6% simple interest for 6 years


Answer by rcmcc(152)

For compounded interest use the equation A=P(1+i)^n and simple interest use the equation A=Prt
First we can eliminate choices with common knowledge b, compounded annually will earn less interest that c. would compounding quarterly.
a)1000(1+.06)^5=1338.23
C)1000(1+(.07/4))^16=1319.93
D)1000(.07)6=1420.00
d) will produce the greatest amount of money.